OpenStudy (anonymous):

If 75 equals a number plus four times the number, find the number. A. –15 B. 30 C. 15 D. –30

OpenStudy (anonymous):

75=n+4(n) is that the equation?

OpenStudy (anonymous):

simplified would be 75=n+4n, so combine like terms. n+4n is 5n. now its 75=5n. divide by 5 on both sides. your answer is15.

OpenStudy (anonymous):

so its c.

OpenStudy (anonymous):